OpenStudy (kawii2004):

Which is the side length of a cube with a volume of 216 in3? A. B. C. D.

OpenStudy (xguardians):

Answr choices?

OpenStudy (xguardians):

Volume = LWH

OpenStudy (xguardians):

Therefore, 216 = x^3 (since all the sides are equal on a square)

jimthompson5910 (jim_thompson5910):

Hint: let's say the cube had a side length of 4 that would mean volume = (length)*(width)*(height) = 4*4*4 = 64 cubic units

OpenStudy (kawii2004):

no not this one one sec

jimthompson5910 (jim_thompson5910):

so IF it said that the volume was 64, then you go in reverse to use the cube root to get 4 however, this is just an example

jimthompson5910 (jim_thompson5910):

hopefully that makes sense @kawii2004 ?